Circular No. 9005

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 X4 = P/2003 K2 (CHRISTENSEN) K. Battams, Naval Research Laboratory, writes that A. Watson found

a comet on SECCHI HI1-B images (via the STEREO-B spacecraft; cf. IAUC 8926, 8930), Battams noting a very faint, short tail (“discovery” observation tabulated below). R. Kracht made measurements of the object’s position and computed a parabolic orbit, from which M. Meyer suggested identity with comet P/2003 K2 (cf. IAUC 8136, 8138, 8145). The indicated correction to the prediction on MPC 56802 is ΔT = −22 days.

2008 UT  α_{2000}  δ_{2000}  Mag.
Dec. 8.239  12^{h}56.^{m}2  −17^{o}06′  10

Astrometric measurements of P/2008 X4 by Battams, with reductions by B. G. Marsden, are published on MPEC 2008-X81, together with the following orbital elements by Marsden — linking the two apparitions (from 227 observations; mean residual 2′′.9) — and an ephemeris.

Epoch = 2003 Mar. 22.0 TT

T = 2003 Apr. 7.8452 TT  ω = 345.^{o}4643
e = 0.828360  Ω = 93.9292 }2000.0
q = 0.548752 AU  i = 10.1398

a = 3.197110 AU n^{o} = 0.1724121 P = 5.72 years Epoch = 2008 Nov. 30.0 TT

T = 2008 Dec. 19.9610 TT  ω = 345.^{o}7643
e = 0.831673  Ω = 93.8928 }2000.0
q = 0.534811 AU  i = 10.2156

a = 3.177213 AU n^{o} = 0.1740342 P = 5.66 years

COMETS 208P/McMILLAN AND 209P/LINEAR Comet P/2008 U1 (McMillan) = P/2000 S7 (cf. IAUCs 8997, 9000) has

been given the permanent designation 208P. Comet P/2008 X2 (LINEAR) = P/2004 CB (cf. IAUC 9002) is assigned the permanent designation 209P.

V5580 SAGITTARII W. Liller, Vina del Mar, Chile, provides a revised position for this

nova, which should replace his position published on IAUC 9004: α = 18^{h}22^{m}01^{s}.45 ± 0^{s}.06, δ = −28^{o}02′38′′.1 ± 3′′.5 (equinox 2000.0).

2008 December 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green